Grade 11Physical Chemistry

find density of a 3.6M sulpuric acid solution that is 29% h2so4 by mass?

sathwika
8 Years agoGrade 11

1 Answer

gaurang gupta
8 Years ago
Molarity of acid = 3.6;
Molar mass of H2SO4 = 98g
Mass of H2SO4 present in 1 L (1000 mL) of acid = 3.6 x 98 g = 352.8 g
Let 'd' be the density of the acid.
Density = Mass/Volume 
Mass of 1000 mL of acid = (d x 1000) g 
Percentage of H2SO4 in acid = 29
352.9 x 100/1000 d = 29
352.9/10 d = 29
290 d = 352.9
Therefore d = 352.9/290= 1.22 g/mL